OpenStudy (calculusxy):

A stop sign — a regular octagon — can be formed from a 12-inch square sheet of metal by making four straight cuts that snip off the corners. How long, to the nearest 0.01 inch, are the sides of the resulting polygon?

OpenStudy (sshayer):

 \[12-2x=x \sqrt{2},2x+x \sqrt{2}=12\] \[x \left( 2+\sqrt{2} \right)=12,x=\frac{ 12 }{ 2+\sqrt{2} }\] rationalize both sides and find x and then 12-2x .